# Eye of the vortex: bound spectra in tunable horizonless rotational analogs

### v2

@@ Abstract
Analog gravity experiments are making remarkable strides in unveiling both the classical and quantum nature of black holes. By harnessing diverse states of matter, contemporary tabletop setups now replicate strong-field phenomena typically confined to the enigmatic regions surrounding black holes. Through these modern gravity simulators, physical processes once considered elusive may finally be brought into experimental reach. In this work, we investigate the bound spectrum of massless scalar excitations propagating within the effective geometry of a rotating acoustic metric. Specifically, we utilize an analog vortex endowed with a tunable parameter that emulates the spacetime of a rotating gravitational background. This model accommodates both the presence of a sonic horizon–characteristic of an acoustic black hole–for non-zero tuning parameters, and its absence when the parameter vanishes, yielding a horizonless, purely rotational vortex flow devoid of radial inflow. We focus on the latter case, where the vortex flow is purely rotational, and compute the spectral properties of the analog system. The resulting bound-state spectrum is found to be qualitatively consistent with that observed in recent experimental realizations of superfluid Helium giant quantum vortices featuring solid or hollow cores. This correspondence suggests that the analog spacetime geometry used here holds significant potential to replicate the phenomenology of cutting-edge laboratory experiments. In doing so, it offers new insight into the intricate landscape of analog black hole spectroscopy and, potentially, the physical topography of bounded, rotating astrophysical environments around black holes.
@@ I Introduction
Concurrently, analog gravity simulators have shown great progress in the field of classical and quantum BH phenomenology [5, 6, 7, 8, 9, 10]. Contemporary technological advances in controlling and manipulating such emulators have enabled the experimental realization of acoustic analogs of gravitational BHs [11]. The past decade witnessed a large variety of laboratory experiments testing a diverse set of aspects of BH physics that led to tentative evidence of analog Hawking radiation [12, 13, 14, 15, 16, 17], superradiance [18, 19] and imprints of QNMs in perturbed analogs [20, 21, 22]. Bridging the gap between analog BH experiments and realistic GWs is now in full bloom. The latest experimental and theoretical developments in these areas illustrate that there is an overabundance of possibilities, even at cosmological scales [23, 24, 25].
@@
Most BH analogs that are built with fluids [33, 34, 35, 36, 37, 38], optical lasers [39, 40, 41, 42, 43], Bose-Einstein condensates [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], superconductors [55, 56, 57] and even quantum circuits [58, 59, 60, 61] have one thing in common; an acoustic horizon. Recent experiments also utilize superfluids [62, 63, 64, 65, 66, 67], such as superfluid Helium. In some particular superfluid-Helium cases, the acoustic horizon can even be absent, i.e. there is no radial inflow of surface excitations. Recently, such laboratory contraption has been constructed and dubbed a giant quantum vortex [68], due to the enormous amount of vortex quanta used to build the vortex [69]. It has been shown that such experiments, when perturbed with surface waves, produce both QNMs and QBSs, regardless of the absence of an event horizon [68, 22]. Utilizing such a circulating vortex without radial flow, and calculating its spectrum is, nonetheless, absent from the current literature.
In this work we revisit an effective metric that describes acoustic excitations on a rotating geometry, with a tuning parameter that can turn the radial inflow on and off. This geometry can describe both the vortex flow experiments involving superfluid Helium without radial flow [68, 22], and rotating acoustic BH analogs with radial flow. Here, we focus on the case of a horizonless acoustic analog, where radial flow is absent, in an attempt to qualitatively replicate the results presented in [68], regarding the QBSs. We find both the co-rotating and counter-rotating QBSs of the rotational horizonless vortex. These are qualitatively compared with the results in [68]. Our analysis matches very well the qualitative results for co-rotating and counter-rotating QBSs, thus we believe that the theoretical setup used here can help into further elucidating the perturbative phenomenology of rotational horizonless vortices, which otherwise is experimentally sensitive to nonlinear dispersion and mechanical noise.
II Rotating acoustic black holes
We begin by considering the acoustic BH solution in Minkowski spacetime obtained by Unruh [11]. Then we discuss a choice for the fluid flow velocity in order to obtain a rotating acoustic BH [5, 7] in cylindrical coordinates, which could also be named laboratory coordinates. By switching off the tuning parameter related to the inflow, we obtain the acoustic metric describing a purely rotational fluid flow.
The fundamental equations of motion for an irrotational fluid are given by
×$\nabla\times\mathbf{v}$ $=$ $0,$ (1)
@@ II Rotating acoustic black holes
with $\vec{v}=v_{0}^{r}\hat{r}+v_{0}^{\theta}\hat{\theta}$. Since we included rotation, the equation $c_{s}^{2}-(v_{0}^{r})^{2}-(v_{0}^{\theta})^{2}=0$ provides a condition for finding the effective ergosphere radius, which in turn are connected with energy extraction from BHs, and superradiance [70]. The radial and angular components of the fluid flow velocity can be written as
@@
III Horizonless vortex fluid flow
When one nullifies the drain, i.e. when $D=0$, then the vortex becomes purely rotational and radial fluid flow is absent. In this case, the background presents no acoustic event horizon, no radial inflow, and hence it describes a horizonless vortex fluid flow; an effective geometry that has been recently constructed in the laboratory with the use of superfluid Helium [68].
@@ A $(2+1)$-dimensional metric describing such a vortex model can be written as
where we have set $z=h_{0}=\mbox{constant}$. In what follows, we will analyze the propagation of scalar excitations at the region of the general vortex spacetime given by Eq. (23), with $D=0$, and calculate the bound spectra. We will adopt an approach that includes the transformation of the equation of motion onto a Mathieu differential equation. We will then use the possible pairs of boundary conditions to obtain bound states (BSs) numerically, compare them with respect to different boundary conditions, and finally try to comprehend if there is a boundary condition that is more physical and closer to the giant quantum vortex experiment [68].
III.1 Analogy between the effective metric used and modern laboratory experiments: an interlude
@@
In particular, the experimental setup that we will try to qualitatively replicate from Refs. [68, 22] has a lot of similarities with the geometry described by Eq. (23). There, they investigate phenomena in the limit of negligible viscosity by using cooled-down superfluid 4He. Specifically, at $T=1.75$ K, at which the experiment takes place, the kinematic viscosity of the superfluid is reduced by a factor of $100$ with respect to water experiments. The key feature of superfluid 4He is the fact that it supports topological defects known as quantum vortices that carry circulation quanta. Each circulation quantum forms an irrotational (zero-curl) flow field in its vicinity. Due to such discretization, a vortex of superfluid 4He can manifest itself only as a giant (multiply quantized) vortex. Such vortices can mimic rigid-body rotation and altogether suggest that an extensive draining vortex of 4He is a feasible candidate for experimental simulations of a quantum-field theory in curved spacetime. An important component of the setup in [68, 22] is a spinning propeller, which establishes a continuous circulation. We can loosely identify the propeller frequency with the parameter $C$ in Eq. (23). Even though the resulting circulation stems from circulation quanta, a record-breaking amount of these quanta have been used ($\sim 4\times 10^{4}$) that allows for the quantization of circulation to be disregarded; the giant quantum vortex is effectively classical. The most significant aspect of the apparatus in [68, 22] is the absence of radial flow of interface fluctuations. This means that the experiment includes pure circulation and no acoustic horizon, as it is the case for the metric in Eq. (23), when $D=0$.
Note that the geometry used here assumes the hydrodynamic limit, which effectively cancels any potential nonlinearities that emerge from dispersion. Not taking into consideration dispersive, nonhydrodynamic, effects practically leads to Unruh’s historical acoustic metric proposal [11] that formed the onset of the interest in hydrodynamic analog models of gravity and the construction of laboratory experiments. Nevertheless, avoiding dispersion in laboratory experiments is extremely difficult since nonlinearities are always part of realistic analog experiments [33, 71, 47, 72, 73, 68, 22]. Even so, simplified BH analog metrics like the one used here, can competently simulate the basic features and dynamics that take place in realistic analog BH experiments.
In this work, we will qualitatively replicate some results that have been obtained experimentally in [68], regarding interface fluctuations in circulating superfluid Helium and the resulting QBSs. We note that our perturbative scheme involves acoustic, non-dispersive, waves which is quite different from the dispersive interface waves of [68]. This is due to the use of a superfluid. Our metric is independent of the fluid’s state or in general the state of matter used to create the vortex. Therefore, it can be applied for other experiments as well, though here we will commence the study by testing the robustness of our analysis by comparing our results with those found in [68].
@@ IV.1 Scalar wave equation
We are interested in the basic features of the effective geometry given by Eq. (23), in particular the propagation of excitations on the surface of the effective fluid flow. In order to perform this analysis, we have to solve the wave equation (9) with appropriate boundary conditions to obtain the QBSs. By substituting Eqs. (23) into Eq. (9), we obtain the following scalar wave equation
$-\frac{\partial^{2}\Psi_{1}}{\partial t^{2}}+\frac{1}{r}\frac{\partial}{\partial r}\biggl{(}r\frac{\partial\Psi_{1}}{\partial r}\biggr{)}+\biggl{(}\frac{1}{r^{2}}-\frac{C^{2}}{r^{4}}\biggr{)}\frac{\partial^{2}\Psi_{1}}{\partial\theta^{2}}-\frac{2C}{r^{2}}\frac{\partial^{2}\Psi_{1}}{\partial t\partial\theta}=0,$ (24)
@@
where $\omega$ is the frequency (energy, in the natural units), $R(r)$ is the radial function, and $m$ is the azimuthal number (a discrete integer running from $\infty$- ∞ to +$\infty$+ ∞). By substituting Eqs. (23)-(25) into Eq. (9), we obtain a radial master equation given by
@@
IV.2 Bound states
Equation (26) can be re-casted into the form of a Mathieu differential equation [74] in order to be solved for its eigenvalues and eigenfunctions. We define a new radial coordinate $x$ such that
$r=\tau\mbox{e}^{ix}\,,\quad\quad x\in\mathbb{R}.$ (27)
@@ IV.2 Bound states
$R^{\prime\prime}(x)+\biggl{(}m^{2}+2Cm\omega-\frac{C^{2}m^{2}\mbox{e}^{-2ix}}{\tau^{2}}-\tau^{2}\omega^{2}\mbox{e}^{-2ix}\biggr{)}R(x)=0,$ (28)
@@
$R^{\prime\prime}(x)+[m(m+2C\omega-2C\omega\cos(2x)]R(x)=0,$ (29)
@@
This transformation can be understood as a mapping from the real variable $x\in\mathbb{R}$ to a complex-valued radial coordinate $r(x)$ lying on a circle of radius $\tau$ in the complex plane. Thus, this transforms Eq. (26) to the following normal form
$R^{\prime\prime}(x)+[a-2q\cos(2x)]R(x)=0.$ (31)
That is the Mathieu equation [75], whose general solution is given by
$R(x)=C_{1}\ \mbox{MathieuC}(a,q,x)+C_{2}\ \mbox{MathieuS}(a,q,x),$ (32)
where $C_{1}$ and $C_{2}$ are constants to be determined, MathieuC and MathieuS are the even and odd Mathieu functions, respectively, and the parameters $a$ and $q$ are given by
@@
To obtain the eigenfrequencies, we need to impose appropriate boundary conditions on the radial solution given by Eq. (32). The boundary value problem related to Mathieu functions will define the physical meaning of the solutions with respect to an “idealized” version of the laboratory experiment we are trying to qualitatively reproduce. In fact, since the Mathieu functions are periodic, the solutions we will obtain from our idealized experiment will be BSs, instead of QBSs. Nevertheless, we will see that to a certain extent, these solutions describe the qualitative aspects of the dynamics of the actual laboratory experiment in [68].
@@
The Mathieu functions become periodic for some characteristic values of $a$ for given $q$. In any half-open interval of length $\pi$ on the real axis, an even Mathieu function with $n$ zeros in the interval $0\leq x<\pi$, $\mbox{MathieuC}_{n}(q,x)$, has characteristic values denoted by $A_{n}(q)$, with $n=0,1,2,\ldots$, while an odd Mathieu function, $\mbox{MathieuS}_{n}(q,x)$, has characteristic values denoted by $B_{n}(q)$, with $n=1,2,\ldots$. These characteristic values satisfy any of the following pairs of boundary conditions [75]:
(I) $\qquad R(0)=R(\pi)=0,\ \mbox{for}\ \mbox{MathieuS}_{n}(q,x)\ \mbox{with period of}\ \pi;$ (35)
(II) $\qquad R^{\prime}(0)=R^{\prime}(\pi)=0,\ \mbox{for}\ \mbox{MathieuC}_{n}(q,x)\ \mbox{with period of}\ \pi.$ (36)
On the other hand, if $R(x)$ is either $\mbox{MathieuS}_{n}(q,x)$ or $\mbox{MathieuC}_{n}(q,x)$, then $R(x)$ and $R(\pi-x)$ satisfy the same differential equation and the same boundary conditions, and must be constant multiples of each other so that $R(x)$ is either an even or an odd function of $\pi/2-x$. This leads to four additional pairs of boundary conditions [75]:
(III) $\qquad R(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuS}_{2n+2}(q,x)\ \mbox{with period of}\ \pi;$ (37)
(IV) $\qquad R(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuS}_{2n+1}(q,x)\ \mbox{with period of}\ 2\pi;$ (38)
(V) $\qquad R^{\prime}(0)=R(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuC}_{2n+1}(q,x)\ \mbox{with period of}\ 2\pi;$ (39)
(VI) $\qquad R^{\prime}(0)=R^{\prime}(\pi/2)=0,\ \mbox{for}\ \mbox{MathieuC}_{2n}(q,x)\ \mbox{with period of}\ \pi.$ (40)
Here, for each $n=0,1,2,\ldots$ there is exactly one characteristic function of each of these four boundary value problems, and $n$ is the number of zeros in the interval $20<x<\pi/$ 2.
@@
Figure 2: Amplitudes of acoustic waves $|R^{(i)}_{mn}(r)|$ corresponding to the $m=1$ mode, with unitary circulation $C=1$. Here, we plot the first three (non-zero) overtones of the BSs, with respect to the radius $r$ ranging from 0 to $\pi$ and frequencies $\omega^{(i)}_{mn}$, for all possible boundary conditions (i=I-VI). We also show the effective potential $V^{(i)}_{{\rm eff},mn}$ with gray curves [the $\omega$-dependent factor, multiplying $R(r)$ in Eq. (26)] for different BS frequency overtones. The solid colored lines correspond to the amplitude of rescaled BSs, which show a characteristic oscillatory pattern, that can be identified with their corresponding BS frequencies (dashed colored lines) outside the potential. The leftmost vertical dashed line corresponds to a reference point at $r=0.8$, while the rightmost one corresponds to $r=\pi$, as seen in the left subfigure of the effective potential, and $r=0.8$ as seen in the right subfigure of the BS amplitudes. Crossing points with the effective potential are marked by gray points.
@@
Figure 3: Comparison between the absolute value of the wave eigenfunctions $|R^{(i)}_{mn}(x)|$, in the $x$-space of Mathieu functions, for all possible boundary conditions (i=I-VI). Here, the azimuthal number is $m=1$ and the circulation is $C=1$, where we vary the overtone number $n$.
From a purely mathematical perspective, we cannot exclude any of the boundary conditions. Thus, on the solution for the radial part of the massless Klein-Gordon equation in a vortex rotational fluid flow, given by Eq. (32), we will impose all six pairs of boundary conditions, so that the parameter $a$ becomes characteristic value of one of the periodic Mathieu functions, i.e.,
$a=B_{N}(q),\ \mbox{for}\ \mbox{MathieuS}_{N}(q,x),$ (41)
or
$a=A_{N}(q),\ \mbox{for}\ \mbox{MathieuC}_{N}(q,x).$ (42)
[…]

### v3

# Vortices without inflow: bound spectra in horizonless rotational analogs

@@ Abstract
Analog gravity experiments are making remarkable strides in unveiling both the classical and quantum nature of black holes. By harnessing diverse states of matter, contemporary tabletop setups now replicate strong-field phenomena typically confined to the enigmatic regions surrounding black holes. Through these modern gravity simulators, physical processes once considered elusive may finally be brought into experimental reach. In this work, we investigate the spectrum of massless scalar excitations propagating within the effective geometry of a rotating acoustic metric. Specifically, we build an analog vortex-like spacetime endowed with a tunable parameter that emulates the geometry of a rotating gravitational background. This model accommodates both the presence of a sonic horizon, characteristic of an acoustic black hole for non-zero tuning parameters, and its absence when the parameter vanishes, yielding a horizonless, purely rotational vortex flow devoid of radial inflow. We focus on the case where the vortex flow is purely rotational. The resulting spectral properties is found to be qualitatively consistent with that observed in recent experimental realizations of giant multiply quantum vortices featuring solid or hollow cores. This correspondence suggests that the analog spacetime used here holds significant potential to replicate, qualitatively, the phenomenology of cutting-edge laboratory experiments. In doing so, it offers new insight into the intricate landscape of analog black-hole spectroscopy and, potentially, the resonant topography of bounded, rotating astrophysical environments around black holes.
@@ I Introduction
Concurrently, analog gravity simulators have shown great progress in the field of classical and quantum BH phenomenology [5, 6, 7, 8, 9, 10]. Contemporary technological advances in controlling and manipulating such emulators have enabled the experimental realization of acoustic analogs of gravitational BHs [11]. The past decade witnessed a large variety of laboratory experiments testing a diverse set of aspects of BH physics that led to tentative evidence of analog Hawking radiation [12, 13, 14, 15, 16, 17], superradiance [18, 19] and imprints of QNMs in perturbed analogs [20, 21]. Bridging the gap between analog BH experiments and realistic GWs is now in full bloom. The latest experimental and theoretical developments in these areas illustrate that there is an overabundance of possibilities [22], even at cosmological scales [23, 24, 25].
@@
Most BH analogs that are built with fluids [33, 34, 35, 36, 37, 38], optical lasers [39, 40, 41, 42, 43], Bose-Einstein condensates [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56], superconductors [57, 58, 59] and even quantum circuits [60, 61, 62, 63] have one thing in common; an acoustic horizon. Recent experiments also utilize superfluids [64, 65, 66, 67, 68, 69], such as superfluid Helium. In some particular superfluid-Helium cases, the acoustic horizon can even be absent, i.e. there is no radial inflow of surface excitations. Recently, such laboratory contraptions have been constructed, dubbed multiply giant quantum vortex [70, 71, 72], due to the enormous amount of vortex quanta used to build the vortex [73]. It has been shown that such experiments, when perturbed with surface waves, produce both QNMs and QBSs, regardless of the absence of an event horizon [70, 71, 72]. Utilizing such a circulating vortex without radial flow, and calculating its spectrum is, nonetheless, absent from the current literature.
In this work we revisit an effective metric that describes acoustic excitations on a rotating geometry, with a tuning parameter that can turn the radial inflow on and off. This geometry can describe both the vortex flow experiments involving superfluid Helium or quantum polaritonic fluids of light without radial flow [70, 71, 72], and rotating acoustic BH analogs with radial flow. Here, we focus on the case of a horizonless acoustic analog, where radial flow is absent, in an attempt to qualitatively replicate the results presented in [70], regarding the QBSs. We find both the co-rotating and counter-rotating QBSs of the rotational horizonless vortex. These are qualitatively compared with the results in [70]. Our analysis matches well the qualitative results for co-rotating and counter-rotating QBSs, thus the theoretical setup used here can help into further elucidate the perturbative phenomenology of rotational horizonless vortices, which otherwise are experimentally sensitive to nonlinear dispersion and mechanical noise [70, 71, 72].
II Rotating acoustic tunable analogs
We consider the acoustic BH solution in Minkowski spacetime obtained by Unruh [11]. Then, we discuss a choice for the fluid flow velocity in order to obtain a rotating acoustic BH [5, 7] in cylindrical coordinates. By switching off the tuning parameter related to the radial inflow, we obtain an acoustic metric describing a purely rotational fluid flow. The fundamental equations of motion for an irrotational fluid are given by
$\nabla\times\mathbf{v}$ $=$ $0,$ (1)
@@ II Rotating acoustic tunable analogs
with $\vec{v}=v_{0}^{r}\hat{r}+v_{0}^{\theta}\hat{\theta}$. Since we included rotation, the equation $c_{s}^{2}-(v_{0}^{r})^{2}-(v_{0}^{\theta})^{2}=0$ provides a condition for the effective ergosphere radius, which in turn is connected with energy extraction from BHs, and superradiance [74]. The radial and angular components of the fluid flow velocity can be written as
@@
III A vortex without radial inflow
When one nullifies the drain, i.e. when $D=0$, then the vortex becomes purely rotational and radial fluid flow is absent. In this case, the background presents no acoustic event horizon, due to the absence of radial inflow, and hence it describes a horizonless vortex fluid flow; an effective geometry that has been recently constructed in the laboratory with the use of superfluid Helium [70, 71] and quantum fluids of light [72].
@@ A $(2+1)$-dimensional metric describing such a vortex model can be written as
where we have set $z=h_{0}=\mbox{constant}$. In what follows, we will analyze the propagation of scalar excitations at the region of the general vortex spacetime given by Eq. (23), with $D=0$, and calculate the bound spectra. We will adopt an approach that includes the transformation of the equation of motion into a modified Mathieu differential equation. We will then obtain the bound states (BSs) numerically, as characteristic values of the modified Mathieu functions, which are intrinsically connected with the standard Mathieu functions [75]. In fact, the BSs are also solutions to the six boundary value problems that the standard Mathieu differential equation allows, thus they are eigenvalues [76, 77, 78]. Finally, we will attempt to comprehend if there is a set of eigenfunctions that are more “physical” and closer to the qualitative behavior of the giant quantum vortex experiment in Ref. [70].
III.1 Analogy between the effective metric used and modern laboratory experiments
@@
In particular, the experimental setup that we will try to qualitatively replicate from Ref. [70] has a lot of similarities with the geometry described by Eq. (23). The key feature of the experiment in [70] is the use of superfluid 4He, that supports topological defects, known as quantum vortices, that carry circulation quanta. Each circulation quantum forms an irrotational (zero-curl) flow field in its vicinity. Due to such discretization, a vortex of superfluid 4He can manifest itself only as a giant multiply quantized vortex. Such vortices can mimic rigid-body rotation and altogether suggest that an extensive draining vortex of 4He is a feasible candidate for experimental simulations of a quantum-field theory in curved spacetime. The setup also features a propeller that gives continuous circulation that can be loosely identified with the parameter $C$ in Eq. (23). The most significant aspect of the apparatus in [70] is the absence of radial flow of interface fluctuations. This means that the experiment includes pure circulation and no acoustic horizon, as it is the case for the metric in Eq. (23).
The geometry used here assumes the hydrodynamic limit, which effectively cancels any potential nonlinearities that emerge from dispersion. Not taking into consideration dispersive, nonhydrodynamic, effects practically leads to Unruh’s historical acoustic metric proposal [11] that formed the onset of the interest in hydrodynamic analog models of gravity and the construction of laboratory experiments. Nevertheless, avoiding dispersion in laboratory experiments is extremely difficult since nonlinearities are always part of realistic analog experiments [33, 79, 48, 80, 81, 70, 71, 72]. Even so, simplified BH analog metrics like the one used here, can competently simulate the basic features and dynamics that take place in realistic analog BH experiments.
Our perturbative scheme involves acoustic, non-dispersive, waves which is quite different from the dispersive interface waves of [70]. This is due to the use of a superfluid. Our metric is independent of the fluid’s state or in general the state of matter used to create the vortex [70, 71, 72]. Therefore, it can be applied for other experiments as well, though here we will commence the study by testing the robustness of our analysis by comparing our results with those in [70].
@@ IV.1 Scalar wave equation
We are interested in the basic features of the effective geometry given by Eq. (23), in particular the propagation of excitations on the surface of the effective fluid flow. In order to perform this analysis, we have to solve the wave equation (9). By substituting Eqs. (23) into Eq. (9), we obtain the following scalar wave equation
$-\frac{\partial^{2}\Psi_{1}}{\partial t^{2}}+\frac{1}{r}\frac{\partial}{\partial r}\biggl{(}r\frac{\partial\Psi_{1}}{\partial r}\biggr{)}+\biggl{(}\frac{1}{r^{2}}-\frac{C^{2}}{r^{4}}\biggr{)}\frac{\partial^{2}\Psi_{1}}{\partial\theta^{2}}$
$-\frac{2C}{r^{2}}\frac{\partial^{2}\Psi_{1}}{\partial t\partial\theta}=0,$ (24)
@@
where $\omega$ is the frequency (energy, in the natural units), $R(r)$ is the radial function, and $m$ is the azimuthal number (a discrete integer running from -$\infty$ to $\infty$). By substituting Eqs. (23)-(25) into Eq. (9), we obtain a radial master equation given by
@@
IV.2 Bound spectra
Equation (26) can be re-casted into the form of a modified Mathieu differential equation [82] in order to be solved for its eigenvalues and eigenfunctions. We define a new radial coordinate $x$ such that
$r=\tau\mbox{e}^{x},\qquad r\in(0,\infty),\qquad x\in(-\infty,\infty).$ (27)
@@ IV.2 Bound spectra
$R^{\prime\prime}(x)-\biggl{(}m^{2}+2Cm\omega-\frac{C^{2}m^{2}\mbox{e}^{-2x}}{\tau^{2}}-\tau^{2}\omega^{2}\mbox{e}^{-2x}\biggr{)}R(x)=0,$ (28)
@@
$R^{\prime\prime}(x)-[m^{2}+2Cm\omega-2Cm\omega\cosh(2x)]R(x)=0,$ (29)
@@
Thus, this transforms Eq. (26) to the following normal form
$R^{\prime\prime}(x)-[a-2q\cosh(2x)]R(x)=0.$ (31)
This is the modified Mathieu equation [76, 77, 75], whose general solution is given by
$R(x)$ $=C_{1}\ \mbox{MathieuC}(a,q,ix)$
$+C_{2}\ \mbox{MathieuS}(a,q,ix),$ (32)
where $C_{1}$ and $C_{2}$ are constants to be determined, MathieuC and MathieuS are given in terms of the even and odd standard Mathieu functions, respectively, and the parameters $a$ and $q$ are given by
@@
Note that the modified Mathieu equation (31) (and its eigenfunctions) differs from the standard Mathieu equation (41) (and its eigenfunctions) only by the input parameter $ix$ in the modified Mathieu functions and $x$ in the standard Mathieu functions. Frequently, Eq. 31) appears in conjunction with Eq. (41), when $a$ has one of the characteristic values $a_{n}(q)$ or $b_{n}(q)$, discussed below. We will restrict ourselves to this case, and obtain the BS frequencies as eigenvalues of the radial solution given by Eq. (32). For a short review on the Mathieu functions, see Appendix A.
The boundary value problems related to the modified Mathieu functions will define the physical meaning of the solutions with respect to an “idealized” version of the laboratory experiment we are trying to qualitatively reproduce. In fact, since the modified Mathieu functions have real eigenvalues, the eigensolutions we will obtain from our idealized experiment will be BSs, instead of QBSs. Nevertheless, we will see that to a certain extent, these eigensolutions describe the qualitative aspects of the dynamics of the actual laboratory experiment in [70].
From a purely mathematical perspective, we cannot exclude any of the six eigenvalues. Thus, on the solution for the radial part of the massless Klein-Gordon equation in a vortex rotational fluid flow, given by Eq. (32), we will impose all six pairs of boundary value conditions, so that the parameter $a$ becomes eigenvalues of one of the six modified Mathieu functions, i.e.,
(I) $\qquad a=b_{n},\ \mbox{for}\ \mbox{Se}_{n}(x);$ (35)
(II) $\qquad a=a_{n},\ \mbox{for}\ \mbox{Ce}_{n}(x);$ (36)
(III) $\qquad a=b_{2n+2},\ \mbox{for}\ \mbox{Gey}_{2n+2}(x);$ (37)
(IV) $\qquad a=b_{2n+1},\ \mbox{for}\ \mbox{Gey}_{2n+1}(x);$ (38)
(V) $\qquad a=a_{2n+1},\ \mbox{for}\ \mbox{Fey}_{2n+1}(x);$ (39)
(VI) $\qquad a=a_{2n},\ \mbox{for}\ \mbox{Fey}_{2n}(x).$ (40)
Here, $\mbox{Se}_{n}(x)=-i\ \mbox{se}_{n}(ix)$ and $\mbox{Ce}_{n}(x)=\mbox{ce}_{n}(ix)$ are the modified Mathieu functions of the first kind. Furthermore, $\mbox{Gey}_{n}(x)$ and $\mbox{Fey}_{n}(x)$ are the modified Mathieu functions of the second kind, which are given in terms of products between Mathieu functions of the first kind and Bessel functions of the second kind [75]. In fact, the parameters $a$ and $q$ are functions of $m$, $C$, $\omega$ and now of $n$. Hence, from the boundary value conditions given by Eqs. (35)-(40) we can find the eigenvalues $\omega^{(i)}_{mn}$, which are the set of BS frequencies with azimuthal number $m$ and overtone number $n$, where $i=(\textrm{I},\ldots,\textrm{VI})$ labels the solution that corresponds to one of the six respective boundary value conditions imposed. To calculate the BS frequencies we solve numerically Eqs. (35)-(40).
@@
Figure 2: Amplitudes of acoustic waves $|R^{(i)}_{mn}(r)|$ corresponding to the $m=1$ mode, with unitary circulation $C=1$. Here, we plot the first three (non-zero) overtones of the BSs, with respect to the radius $r$ ranging from 0 to $\pi$ and frequencies $\omega^{(i)}_{mn}$, for all possible boundary conditions $i=(\textrm{I},\ldots,\textrm{VI})$. We also show the effective potential $V^{(i)}_{{\rm eff},mn}$ with gray curves [the $\omega$-dependent factor, multiplying $R(r)$ in Eq. (26)] for different BS frequency overtones. The solid colored lines correspond to the amplitude of rescaled BSs, which show a characteristic oscillatory pattern, that can be identified with their corresponding BS frequencies (dashed colored lines) outside the potential. The leftmost vertical dashed line corresponds to a reference point at $r=0.8$, while the rightmost one corresponds to $r=\pi$, as seen in the left subfigure of the effective potential, and $r=0.8$ as seen in the right subfigure of the wavefunction amplitudes. Crossing points with the effective potential are marked by gray points.
@@
Figure 3: Comparison between the absolute value of the first three (non-zero) wave eigenfunctions $|R^{(i)}_{mn}(x)|$, in the $x$-space of Mathieu functions, for all possible boundary conditions $i=(\textrm{I},\ldots,\textrm{VI})$. Here, the azimuthal number is $m=1$ and the circulation is $C=1$, where we vary the overtone number $n$.
From a physical point of view, the boundary value problems imposed to the modified Mathieu equation (32) are pivotal to the resulting BS wavefunctions we are trying to reproduce. In the radial context of the problem, and knowing that the hyperbolic cosine will introduce blow-up behavior at $\pm\infty$, we need to apply boundary conditions that lead to regular (square-integrable) solutions. The modified Mathieu equation’s solutions grow or decay exponentially at large $|x|$. So, without constraints, one cannot get physically acceptable BSs. Typical boundary conditions for our problem should normalize the solutions of the modified Mathieu equation in order to obtain well-behaved BSs. These are quite straightforward, and can be found in many contexts such as vibrating systems, electromagnetism, elastic membranes and acoustics [76, 77]. For the problem in hand, we are studying fluctuations on a metric that describes a rotational acoustic analog without inflow. The first restriction comes from the absence of inflow of perturbations at the origin of the vortex $x=0$. We impose either a Dirichlet boundary condition, i.e., $R(0)=0$, or a Neumann boundary, i.e., $R^{\prime}(0)=0$ since solutions must remain finite or regular, there [76, 77]. At infinity, we can either impose purely outgoing waves, which lead to QBS solution that decay as they approach infinity, or impose, again, Dirichlet or Neumann boundary conditions, that lead to BSs. In accord to the conjugation between the modified and the standard Mathieu equations, we shall impose the only mathematically-allowed boundary conditions, shown in Appendix A, Eqs. (43)-(48), that are connected with those in Eqs. (35)-(40), i.e. Dirichlet or Neumann depending on the case. These boundaries seems to satisfy the main principle of BS solutions and regularity [^1] This is analogous to the choice of modified Bessel functions in classical radial problems, where one selects the decaying solutions, at the origin and outer boundary, for BSs..